OpenStudy (anonymous):

1-2n over 3 > or equal to 7

OpenStudy (anonymous):

\[\frac{ 1-2n }{ 3 }\ge 7\]

Parth (parthkohli):

A good start is multiplying both sides by 3.

Parth (parthkohli):

\[1 - 2n \ge 7 \times 3 \]

OpenStudy (anonymous):

okay. what would the next step be?? im getting it

Parth (parthkohli):

The next step is to subtract 1 from both sides.\[-2n \ge 21 - 1 \]